Two fair dice, one blue and one red, are tossed, and the up face on each die is recorded. Define the following events:
E:{ A 1 on the blue die }
F:{ The sum of the numbers is even }
Find the following probabilities:
(a) P(E)=
(b) P(F)=
(c) P(E∩F)=

Found 2 solutions by math_helper, ikleyn:
Answer by math_helper(2461)




P(E) = 1/6, as the up face on the blue die is independent of the red die.

P(F) = 1/2, as 1/2 of all the outcomes are even sums.


P(E∩F) = 1/12, 3 out of 36 possible outcomes fall in the E∩F case.


Answer by ikleyn(52877)

.

If you want to learn BETTER this subject and this class of problems, look into and read my lesson
    - Rolling a pair of fair dice
in this site.


Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic  "Solved problems on Probability".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.